A person pushes a 10 kg cart a distance of 20 meters by exerting a 60 Newton horizontal force. The frictional resistance force is 50 Newtons. How much work is done by each force acting ont he cart? How much kinetic energy does the cart have at the end of the 20 meters if it started from rest:

Grade:12

1 Answers

vikas askiitian expert
13 years ago

work done by friction =50*s =-500J (here force is opposite to displacement)

 workdone by man =60*s=600J      (force is in the direction of displacement)

 net wordone =work by friction + work by man =100J

 kinetic energy change=total work=100J